# Diagonal comparison of Ample $\mathrm{C}^{\ast}$-diagonals

### v1

@@ 4. Hereditary conditional expectations
Now let $\psi\in\mathrm{PS}(D_{2^{\infty}})$ and consider the pure state $\theta\coloneqq\mathrm{ev}_{\infty}\otimes\psi\in\mathrm{PS}(D)$. Clearly $\theta$ will fail to extend uniquely on $B$, since if $\omega\in\beta\mathbb{N}\setminus\mathbb{N}$ is any free ultrafilter, the state $(b_{j})_{j\geq 1}\mapsto\lim_{j\to\omega}\psi(E(b_{j}))$ will extend $\theta$ on $B$. Nevertheless, since $A=\mathrm{C}^{\ast}(D,p)$, if we show that any extension of $\theta$ vanishes at $p$, then a straightforward multiplicative domain argument shows that all extensions of $\theta$ agree on $A$. Since $\psi\in\mathrm{PS}(D_{2^{\infty}})$ and the spectrum of $D_{2^{\infty}}$ is the Cantor space $\{1,2\}^{\mathbb{N}}$, there is a point $x=(x_{j})_{j\geq 1}\in\{1,2\}^{\mathbb{N}}$ such that $\psi$ corresponds to evaluation at $x$. For $k\in\mathbb{N}$ set $\bar{x}_{k}:=(x_{j})_{j=1}^{k}\in\{1,2\}^{k}$ and, for a $k$-tuple $\bar{y}:=(y_{j})_{j=1}^{k}\in\{1,2\}^{k}$ set $e_{\bar{y}}:=e_{y_{1},y_{1}}\otimes\dots\otimes e_{y_{k},y_{k}}\in D_{2}^{\otimes k}\subset D_{2^{\infty}}$. Now let $e_{k}:=(0,\dots,0,e_{\bar{x}_{k}},e_{\bar{x}_{k}},\dots)\in D$, where the segment of zeroes in the beginning of $e_{k}$ has length $k$. Clearly $\theta(e_{k})=1$ for all $k\in\mathbb{N}$, so each $e_{k}$ lies in the multiplicative domain of $\theta$, hence $\theta(p)=\theta(e_{k}pe_{k})$ for all $k\in\mathbb{N}$. We have that $\|e_{k}pe_{k}\|=\sup_{n\geq k+1}\|e_{\bar{x}_{k}}p_{n}e_{\bar{x}_{k}}\|_{M_{2}^{\otimes n}}$, and for $n\geq k+1$ we have
@@
Note that $(D\subset A)$ is non-regular, since $\mathcal{N}_{A}(D)=D$. Also, $A$ is an AF algebra, which we prove by showing that $A$ is locally AF. Since $A$ is a generated by $D\cup\{p\}$, it suffices to show that if $\Omega$ is a finite subset of $D\cup\{p\}$ and $\varepsilon>0$, then there exists a finite-dimensional sub-$\mathrm{C}^{\ast}$-algebra $F\subset A$ such that $\mathrm{dist}(a,F)<\varepsilon$ for all $a\in\Omega$.

### v2

@@ 4. Hereditary conditional expectations
Now let $\psi\in\mathrm{PS}(D_{2^{\infty}})$ and consider the pure state $\theta\coloneqq\mathrm{ev}_{\infty}\otimes\psi\in\mathrm{PS}(D)$. Clearly $\theta$ will fail to extend uniquely on $B$, since if $\omega\in\beta\mathbb{N}\setminus\mathbb{N}$ is any free ultrafilter, the state $(b_{j})_{j\geq 1}\mapsto\lim_{j\to\omega}\psi(E(b_{j}))$ will extend $\theta$ on $B$. Nevertheless, since $A=\mathrm{C}^{\ast}(D,p)$, if we show that any extension $\bar{\theta}$ of $\theta$ on $A$ vanishes at $p$, then since $p$ is a projection we will have that $0=|\bar{\theta}(p)|^{2}=\bar{\theta}(p^{2})$, and so $p$ will lie in the multiplicative domain of $\bar{\theta}$. As $D$ also lies in the multiplicative domain of $\bar{\theta}$ since $\theta\in\mathrm{PS}(D)$, we have that $\bar{\theta}$ vanishes on any product of elements from $D\cup\{p\}$ whenever at least one term $p$ appears. Since $A=\mathrm{C}^{\ast}(D,p)$, this shows that the extension of $\theta$ is uniquely determined on $A$. Now since $\psi\in\mathrm{PS}(D_{2^{\infty}})$ and the spectrum of $D_{2^{\infty}}$ is the Cantor space $\{1,2\}^{\mathbb{N}}$, there is a point $x=(x_{j})_{j\geq 1}\in\{1,2\}^{\mathbb{N}}$ such that $\psi$ corresponds to evaluation at $x$. For $k\in\mathbb{N}$ set $\bar{x}_{k}:=(x_{j})_{j=1}^{k}\in\{1,2\}^{k}$ and, for a $k$-tuple $\bar{y}:=(y_{j})_{j=1}^{k}\in\{1,2\}^{k}$ set $e_{\bar{y}}:=e_{y_{1},y_{1}}\otimes\dots\otimes e_{y_{k},y_{k}}\in D_{2}^{\otimes k}\subset D_{2^{\infty}}$. Now let $e_{k}:=(0,\dots,0,e_{\bar{x}_{k}},e_{\bar{x}_{k}},\dots)\in D$, where the segment of zeroes in the beginning of $e_{k}$ has length $k$. Clearly $\theta(e_{k})=1$ for all $k\in\mathbb{N}$, so each $e_{k}$ lies in the multiplicative domain of $\theta$, hence $\theta(p)=\theta(e_{k}pe_{k})$ for all $k\in\mathbb{N}$. We have that $\|e_{k}pe_{k}\|=\sup_{n\geq k+1}\|e_{\bar{x}_{k}}p_{n}e_{\bar{x}_{k}}\|_{M_{2}^{\otimes n}}$, and for $n\geq k+1$ we have
@@
Note that $(D\subset A)$ is non-regular, since $\mathcal{N}_{A}(D)=D$. Also, $A$ is an AF algebra, which we prove by showing that $A$ is locally AF. Since $A$ is generated by $D\cup\{p\}$, it suffices to show that if $\Omega$ is a finite subset of $D\cup\{p\}$ and $\varepsilon>0$, then there exists a finite-dimensional sub-$\mathrm{C}^{\ast}$-algebra $F\subset A$ such that $\mathrm{dist}(a,F)<\varepsilon$ for all $a\in\Omega$.